# Atomic Fact Decomposition Helps Attributed Question Answering

### v1

@@ 5.1. Evaluation Setups
Table 2. The ablation experiment results on NQ, StrategyQA, and ExpertQA using GPT-3.5. “w/o” represents removal for the mentioned module. The ablation study on Mintaka is shown in Appendix C.2. We mark the better results in bolded.
Methods NQ StragegyQA ExpertQA
[…]
w/o edit 0.648 0.715 - - - 0.442 0.53 - - - 0.394 0.417 - - -
w/o atomic 0.451 0.619 0.9 0.734 0.608 0.225 0.331 0.918 0.487 0.362 0.168 0.308 0.943 0.464 0.285
w/o molecular 0.713 0.785 0.585 0.670 0.643 0.514 0.609 0.598 0.603 0.553 0.425 0.449 0.561 0.499 0.484
w/o re-retrieval 0.609 0.677 0.911 0.777 0.730 0.437 0.533 0.927 0.677 0.594 0.388 0.421 0.937 0.581 0.548
ARE 0.69 0.737 0.91 0.814 0.772 0.463 0.559 0.899 0.689 0.611 0.412 0.438 0.917 0.593 0.569
5.2. Experimental Results and Analysis
The experimental results are displayed in Table 1, and the experimental analyses are listed as follows:
(1) ARE demonstrates significant improvements in both molecular-level attribution evaluation metrics and intent preservation across four datasets for all LLMs. ARE achieves improvements of 68%, 71.6%, 52.9%, and 42.9% over the $Ave-\mathrm{F1_{PP}}$ metric across all datasets. Additionally, ARE also achieves the best performance in the $Ave-\mathrm{F1_{RP}}$ metric, with improvements of 7.3%, 9.4%, 9.2%, and 7.3% on GPT-3.5, Llama3-70B, and Llama2-70B [^8] We use GPT-3.5-turbo-1106 for reproducibility. The Llama series models can be downloaded from https://huggingface.co/meta-llama/Llama-2-70b-chat and https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct., respectively, demonstrating the superior effectiveness and generalization capabilities of our proposed ARE.
@@ 5.2. Experimental Results and Analysis
(2) For attribution ability, ARE outperforms other baselines on $Attr_{r}$ and $Attr_{p}$. Especially in $Attr_{p}$, all existing methods are far below ARE, this is because $Attr_{p}$ is a strict metric that evaluates the completeness of evidence. As for Pres, although the CCVER gets a higher $Pres$ in GPT-3.5 and Llama3-70B, it performed poorly on Llama-2-70B. In contrast, ARE achieved the most stable intent preservation across all LLMs, even with the less effective Llama2-70B model, which struggles with intent preservation, particularly when using the RARR method. These results demonstrate that the atomic fact-base ARE can maximize the attribution score while preserving the original intent.
@@
(3) The sub-question generation methods perform similarly on NQ, Mintaka, and StrategyQA, but CCVER significantly underperforms compared to RARR on ExpertQA. This highlights the limitations of sub-question generation, which require carefully designed prompts. Conversely, ARE shows greater robustness and adaptability across different datasets. Further experimental results are available in Appendix D.
Figure 4. The performance of the fact decomposition LLM at the molecular level and the atomic level for different number of iterations.
5.3. Ablation Study
@@
In order to investigate the impact of key modules on experimental performance, we conduct a series of ablation experiments, and the corresponding results are presented in Table 2. Specifically, “w/o edit” means removing the editing module; “w/o atomic” represent removing atomic-level facts; “w/o molecular” represents removing molecular-level clause; “w/o re-retrieval” means removing the status of “irrelevant” in evidence verifier.
@@
(1) In the “$w/o$ edit” setting, attribution scores decline significantly. For example, $Attr_{r}$ decreased by 4.2%, while $Attr_{p}$ dropped by 2.2% in NQ. These results underscore the importance of atomic fact editing, as modifying hallucinated content is crucial for further improving attribution accuracy.
@@
(2) The results of the “w/o atomic” setting show the atomic facts play an important role in attribution scores. Specifically, ARE obtains 23.9% in $Attr_{r}$ and 11.8% in $Attr_{p}$ improvements on the NQ dataset. This phenomenon indicate that atomic facts have higher certainty and lower entropy, making it possible to retrieve relevant evidence more effectively.
@@
(3) Although the “w/o molecular” setting performs better on $Attr_{r}$ and $Attr_{p}$, it shows poorer performance on the $pres$ metric. This may be attributed to the removal of molecular facts, which causes atomic facts to lose their correspondence with them, preventing accurate backtracking to their original positions for editing. Consequently, this significantly increases the risk of altering the original intent.
@@
(4) The performance of “w/o re-retrieval” shows that the necessity of the “irrelevant” status in evidence verifier module. For example, the $Attr_{r}$ decreased by 8.1% and 6% in $Attr_{p}$ in the NQ dataset. This phenomenon demonstrates that expanding facts when they have irrelevant evidence and then re-retrieving evidence can effectively enhance the attribution capability.
@@
Table 3. Comparison of the proposed atomic fact-based retrieval with previous sub-question-based retrieval performance.
@@
$Attr_{r}$ $Attr_{p}$ $Attr_{r}$ $Attr_{p}$ $Attr_{r}$ $Attr_{p}$
@@
5.4. The Impact of Fact Decomposition LLM with Different Iterations
To find the best performance of molecular-to-atomic fact decomposition LLM and explore how iterations affect fact decomposition performance, we perform a comparative experiment, and the results are illustrated in Figure 4. Specificlly, we assess its performance at both the molecular and atomic levels, focusing on two key aspects: consistency and correctness. Consistency $n_{c}$ measures how closely the number of decomposed sentences aligns with that of gold sentences. Correctness $d_{correct}$ evaluates whether the decomposed sentences preserve the original sentence meaning.
We test on a well-constructed evaluation dataset mentioned in 4.1 to select the decomposition model used in ARE. As shown in Figure 4, the performance of $n_{c}$ at the molecular-level initially increase and then decreasing as the number of epochs increase. At the atomic-level, $n_{c}$ shows an overall upward trend, stabilizing at epoch 4. For $d_{correct}$, performance at molecular-level also increase and then decrease as the number of epochs increase, peaking at epoch 3 with 0.972. However, at the atomic-level, it first decreases and then increases.
In terms of consistency, the performance at the 3rd and 4th epochs is similar. However, for correctness, the score of the atomic-level at epoch 3 is the highest, and meanwhile the score of the molecular-level at epoch 3 and epoch 4 has little difference. Therefore, we select the LLM trained with 3 epochs. Details of evaluation process can be found in Appendix D.5.
Table 4. Comparing the revised results with different systems. Red indicates hallucinated content that conflicts from the evidence or fact. Blue marks the content that, after revision, has been corrected to align with the evidence. Orange indicates it is still not correct after the revision.
@@ 5.4. The Impact of Fact Decomposition LLM with Different Iterations
5.5. Hypothesis Proof Experiment
As illustrated in Table 3, ARE shows the improvements of 4.6% on NQ, 8.3% on StrategyQA, and 9.9% on ExpertQA based on the metric of $Attr_{r}$. This supports the hypothesis that atomic facts are easier to retrieve as relevant evidence. Furthermore, DRQA, using the entire answer and question for retrieval, struggles in cross-document retrieval scenarios. Methods involving generated sub-questions fail to obtain more evidence due to the poor quality of generated questions.
CCVER and RARR both generate sub-questions and then use these sub-questions for retrieval. Their performance is similar on NQ and ExpertQA, but on StrategyQA, CCVER outperforms RARR. This is because CCVER generates “Yes or No” type questions, which match the type of the StrategyQA dataset, allowing for better evidence retrieval. However, both perform worse than ARE, which is based on atomic fact retrieval.
5.6. Visualization results of attribution scores and preservation
To provide a more intuitive comparison between existing methods and our proposed ARE, regarding the simultaneous achievement of maximum attribution scores and preservation of original intent, we conducted the visualization experiments illustrated in Figure 5.
As shown in Figure 5, ARE is positioned on the right side of the dashed line, which indicates that ARE has achieved effective editing. The contour lines indicate that the proposed ARE achieves the best performance in both the F1RP (F1RP both considers Attrr and $Pres$) and F1PP (F1RP simultaneously considers Attrp and $Pres$) metrics after editing. These phenomena demonstrates the robustness of the proposed ARE in maximizing the attribution scores and preserving the original intent. The completed results can be found in Appendix D.4.
5.7. Case Study
@@
When comparing the revised results of EFEC, RARR, and ARE in Table 4, clearly highlights its advantages, as well as the shortcomings of RARR and EFEC. ARE performs minimal edits by focusing on atomic facts, successfully preserving the original intent while improving factuality and attribution. In contrast, RARR often makes substantial modifications, which can make it challenging to preserve the original intent of the content. EFEC performs even less satisfactorily. It makes significant changes to the original content, omitting entire sentences. Such drastic revisions can result in a final product that significantly deviates from the original content and intent.
6. Conclusion

### v2

@@ V-A Evaluation Setups
TABLE III: The ablation experiment results on NQ, StrategyQA, Mintaka and ExpertQA using GPT-3.5. “w/o” represents removal for the mentioned module. We mark the better results in bolded.
Method NQ StragegyQA
$Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$
w/o edit 0.648 0.715 - - - 0.442 0.53 - - -
w/o atomic level 0.451 0.619 0.9 0.734 0.608 0.225 0.331 0.918 0.487 0.362
w/o molecular level 0.713 0.785 0.585 0.670 0.643 0.514 0.609 0.598 0.603 0.553
w/o re-retrieval 0.609 0.677 0.911 0.777 0.730 0.437 0.533 0.927 0.677 0.594
ARE 0.69 0.737 0.91 0.814 0.772 0.463 0.559 0.899 0.689 0.611
Mintaka ExpertQA
w/o edit 0.672 0.768 - - - 0.404 0.417 - - -
w/o atomic level 0.477 0.651 0.928 0.765 0.630 0.168 0.308 0.943 0.464 0.285
w/o molecular level 0.725 0.832 0.548 0.661 0.624 0.425 0.449 0.561 0.499 0.484
w/o re-retrieval 0.68 0.767 0.936 0.843 0.788 0.388 0.421 0.937 0.581 0.548
ARE 0.716 0.807 0.914 0.857 0.803 0.412 0.438 0.917 0.593 0.569
Figure 4: The visualization experiments on NQ and ExpertQA datasets. The dashed line represents the highest attribution scores of all methods before editing. The points represent the performance of various methods after editing. The contours display level curves for $\mathrm{F1_{RP}}$ (left) and $\mathrm{F1_{PP}}$ (right), where points on the same contour share the same value. The closer a point is to the upper right corner, the better its performance is represented.
V-B Experimental Results and Analysis
The experimental results are displayed in Table I and Table II, and the experimental analyses are listed as follows:
(1) ARE demonstrates substantial improvements in both molecular-level attribution evaluation metrics and intent preservation across four datasets using various LLMs. For instance, on Mintaka with API-based LLMs, ARE improves Ave-$\mathrm{F1_{PP}}$ by 72.4% and Ave-$\mathrm{F1_{RP}}$ by 1.6% (Table I). With open-source LLMs (Table II), ARE also achieves state-of-the-art performance—for example, in NQ, Ave-$\mathrm{F1_{PP}}$ and Ave-$\mathrm{F1_{RP}}$ improve by 9.5% and 68.3%, respectively—demonstrating the strong effectiveness and generalizability of our approach. The relatively smaller improvements observed on ExpertQA can be attributed to: (1) its specialized questions that often require expert-level domain knowledge beyond standard open knowledge bases, and (2) its significantly longer answers, averaging over 20 tokens more than other datasets.
@@ V-B Experimental Results and Analysis
(2) For attribution ability, ARE outperforms other baselines in $Attr_{r}$ and $Attr_{p}$. Especially in $Attr_{p}$, all existing methods are far below ARE. This is because CCVER and RARR retrieve a larger quantity of evidence, which is typically fragmented. Although fragmented evidence can be beneficial for $Attr_{r}$ due to its focus on entailment probability, $Attr_{p}$ is a stricter metric that evaluates the completeness of the evidence. These evidence cannot support the entire sentence, resulting in poor performance in $Attr_{p}$. As for Pres, although the CCVER gets a higher Pres in GPT-3.5 and Llama3-70B, it performed poorly on Llama2-70B. In contrast, ARE achieved the most stable Pres across all LLMs, even with the less effective Llama2-70B, particularly when using the RARR method. These results demonstrate that ARE can maximize the attribution score while preserving the original intent.
@@
(3) The sub-question generation methods perform similarly on NQ, Mintaka and StrategyQA, but CCVER significantly underperforms compared to RARR on ExpertQA. This highlights the limitations of sub-question generation, which requires carefully designed prompts. Conversely, ARE shows greater robustness and adaptability across different datasets.
V-C Ablation Study
@@
In order to investigate the impact of key modules on experimental performance, we conduct a series of ablation experiments, and the corresponding results are presented in Table III. Specifically, “w/o edit” means removing the editing module; “w/o atomic” represents removing atomic-level facts; “w/o molecular” represents removing molecular-level clause; “w/o re-retrieval” means removing the status of “irrelevant” in evidence verifier.
@@
(1) In the “w/o edit” setting, attribution scores decline significantly. For example, $Attr_{r}$ decreased by 4.2%, while $Attr_{p}$ dropped by 2.2% on NQ. These results underscore the importance of atomic fact editing, as modifying hallucinated content is crucial for further improving the accuracy of attribution.
@@
(2) The results of the “w/o atomic” setting show that atomic facts play an important role in the attribution scores. Specifically, ARE obtains improvements of 23. 9% in $Attr_{r}$ and 11. 8% in $Attr_{p}$ on NQ dataset. This phenomenon indicates that atomic facts have higher certainty and lower entropy, making it possible to retrieve relevant evidence more effectively.
@@
(3) Although the “w/o molecular” setting performs better on $Attr_{r}$ and $Attr_{p}$, it shows poorer performance on the Pres metric (decreased by 32.5% on NQ, 30.1% on StrategyQA, 35.6% on ExpertQA). This may be attributed to the removal of molecular facts, which causes atomic facts to lose their correspondence with them, preventing accurate backtracking to their original positions for editing. Consequently, this significantly increases the risk of altering the original intent.
@@
(4) The performance of “w/o re-retrieval” shows that the necessity of the “irrelevant” status in evidence verifier module. For example, $Attr_{r}$ decreased by 8.1% and 6% in $Attr_{p}$ in the NQ dataset. This phenomenon demonstrates that expanding facts when they have irrelevant evidence and then re-retrieving evidence can effectively enhance the attribution capability.
TABLE IV: Comparison with two prompt-based variants of ARE on Mintaka dataset using GPT-3.5.
Methods $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$
CCVER 0.630 0.069 0.937 0.129 0.753
RARR 0.646 0.06 0.850 0.112 0.734
Prompt-based two-stage ARE 0.688 0.77 0.858 0.730 0.764
Prompt-based ARE 0.682 0.768 0.774 0.723 0.725
ARE 0.716 0.807 0.914 0.857 0.803
V-D Visualization Results of Attribution Scores and Preservation
To provide a more intuitive comparison between existing methods and our proposed ARE, regarding the simultaneous achievement of maximum attribution scores and preservation of original intent, we conducted the visualization experiments illustrated in Figure 4. ARE is located on the right side of the dashed line, indicating that ARE has achieved effective editing. The contour lines indicate that the proposed ARE achieves the best performance in both the F1RP (F1RP both considers Attrr and Pres) and F1PP (F1RP simultaneously considers Attrp and Pres) metrics after editing. These phenomena demonstrate the robustness of the proposed ARE in maximizing the attribution scores and preserving the original intent.
V-E Effectiveness of Instruction-tuning
To demonstrate the effectiveness of our proposed instruction-tuned fact decomposition LLM, we further present two variants of ARE: 1) Prompt-based two-stage ARE: This method first prompts LLMs to generate molecular clauses and then uses them to generate atomic facts using prompts. 2) Prompt-based ARE: This method only replaces the instruction-tuned fact decomposition LLM with the prompt.
As illustrated in Table IV, ARE significantly outperforms two variants of ARE, the performance decreased by about 12% and 4% in F1PP and F1RP, respectively. Prompt-based ARE fails to capture the complex structure of long-form answers, and prompt-based two-stage ARE still performs less than ARE, although it simplifies the process of generating atomic facts by first generating molecular clauses. Furthermore, ARE gives better performance because it utilizes a fact decomposition LLM tuned on a well-constructed dataset that enables ARE to better capture the complex structure of long-form answers.
V-F Experiments of Aggregating Evidence from Sub-questions
The sub-question generation method (such as RARR) lacks a clear correspondence between questions and clauses, thus failing to provide complete evidence for molecular clauses. To demonstrate this, we employ our Molecular-to-Atomic fact decomposition LLM to first decompose the long-form answer into molecular clauses and then generate sub-questions for every molecular clause using GPT-3.5. Finally, the evidences retrieved through the sub-questions are aggregated according to the molecular clause. As shown in Table V, the performance in $Attr_{p}$ is obviously improved (from 0.109 to 0.766 on NQ) by aggregating the evidence according to the molecular clauses (the method of “RARR w/ aggregate”), showing that $Attr_{p}$ can evaluate the completeness. However, $Attr_{r}$ decreases, indicating that RARR retrieves irrelevant evidence that makes the aggregated evidence noisier.
TABLE V: The experimental results of aggregating evidence retrieved from sub-questions into the corresponding molecular clauses.
Method NQ StrategyQA ExpertQA
$Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$
RARR $w/$ aggregate 0.499 0.646 0.942 0.766 0.653 0.336 0.53 0.948 0.680 0.496 0.276 0.407 0.959 0.571 0.428
V-G The Performance of Fact Decomposition and Quality Analysis of Constructed Dataset
To find the best performance of molecular-to-atomic fact decomposition LLM and explore how iterations affect fact decomposition performance, we perform a comparative experiment, and the results are illustrated in Figure 5. Specifically, we assess its performance at both the molecular and atomic levels, focusing on two key aspects: consistency and correctness. Consistency $n_{c}$ measures how closely the number of decomposed sentences aligns with that of gold sentences. Correctness $d_{correct}$ evaluates whether the decomposed sentences preserve the original sentence meaning.
Figure 5: The performance of the fact decomposition LLM at the molecular level and the atomic level for different number of iterations.
We test on a well-constructed evaluation dataset mentioned in IV-A to select the decomposition model used in ARE. As shown in Figure 5, the performance of $n_{c}$ at the molecular-level initially increases and then decreases as the number of epochs increases. At the atomic-level, $n_{c}$ shows an overall upward trend, stabilizing at epoch 4. For $d_{correct}$, performance at molecular-level also increases and then decreases as the number of epochs increases, peaking at epoch 3 with 0.972. However, at the atomic-level, it first decreases and then increases.
In terms of consistency, the performance at the 3rd and 4th epochs is similar. However, for correctness, the score of the atomic-level at epoch 3 is the highest, and meanwhile the score of the molecular-level at epoch 3 and epoch 4 has little difference. Therefore, we select the LLM trained with 3 epochs.
Although the fact decomposition LLM has achieved satisfactory performance, misaligned decomposition scenarios still exist. Since ARE is a multi-stage pipeline architecture, such misaligned decomposition may lead to error propagation. This issue represents one of the primary challenges in our system. Nevertheless, our multi-stage pipeline design still demonstrates substantial overall gains, and the benefits significantly outweigh the drawbacks introduced by occasional decomposition errors. As shown in Table III, the performance has decreased when removing the editing (“w/o edit”) or re-retrieval (“w/o re-retrieval”) stage respectively, which demonstrates that ARE can address the error propagation.
TABLE VI: The quality report of the generated instruct-tuning dataset
   Dataset    Human    LLM-as-a-Judge
   Train (500)    0.934    0.965
   Test (200)    0.946    0.972
   average    0.940    0.969
We evaluated the quality of the generated instruct-tuning dataset (randomly selected 500 and 200 from train and test respectively) using both human evaluation and LLM-as-a-Judge (Llama-3.1-70B-Instruct). The accuracy scores for LLM generated long-form text, molecular clauses, and atomic facts were 0.94 (human) and 0.969 (LLM), as shown in Table VI.
@@ V-G The Performance of Fact Decomposition and Quality Analysis of Constructed Dataset
TABLE VII: Comparison of the proposed atomic fact-based retrieval with previous sub-question-based retrieval performance.
@@
$Attr_{r}$ $_{p} AttrrAtt$$_{r}$ $_{p} AttrrAtt$$_{r}$ $Attr_{p}$
@@
   Methods
   Latency $\downarrow$
   (sample/s)
   Interaction $\downarrow$
   (times/sample)
   RARR    55    7.1
   CCVER    87    9
   ARE    41    4.7
   Consumed tokens $\downarrow$
   (tokens/sample)
   Retrieval $\downarrow$
   (times/sample)
   RARR    7371    5.93
   CCVER    9711    7
   ARE    5340    4.77
V-H Hypothesis Proof Experiment
As illustrated in Table VII, ARE shows the improvements of 4.6% on NQ, 8.3% on StrategyQA, and 9.9% on ExpertQA based on the metric of $Attr_{r}$. This supports the hypothesis that atomic facts are easier to retrieve as relevant evidence. Furthermore, DRQA, using the entire answer and question for retrieval, struggles in cross-document retrieval scenarios. Methods involving generated sub-questions fail to obtain more evidence due to the poor quality of generated questions.
CCVER and RARR both generate sub-questions and then use them for retrieval. Their performance is similar on NQ and ExpertQA, but on StrategyQA, CCVER outperforms RARR. This is because CCVER generates “Yes or No” type questions, which match the type of the StrategyQA dataset, allowing for better evidence retrieval. However, both perform worse than ARE, which is based on atomic fact retrieval.
V-I Inference Efficiency Analysis
To compare the inference efficiency between ARE and other baselines, we perform controlled inferences CCVER, RARR and ARE on 4 NVIDIA A800 GPUs using VLLM [^9] https://github.com/vllm-project/vllm. We measure average latency, interaction times with LLMs, and token consumption. As shown in Table VIII, ARE achieves a shorter average inference time per sample compared to RARR. The average interaction times with the LLMs are 4.7, and the average tokens consumed are 5340, both of which outperform the baselines. This is because the number of atomic facts generated by instruction-tuned decomposition LLM is less, but they are more relevant to the original answer than RARR and CCVER, leading to fewer retrievals (4.77 times for ARE, 5.93 for RARR and 7 for CCVER).
TABLE IX: Comparing the revised results with different systems. Red indicates hallucinated content that conflicts from the evidence or fact. Blue marks the content that, after revision, has been corrected to align with the evidence. Orange indicates it is still not correct after the revision.
@@ V-I Inference Efficiency Analysis
TABLE X: Error case in ARE. Red indicates the hallucinated content. Blue marks as the correct content. Orange represents the invalid evidence. Green is the specialized word.
Atomic Fact The Euro is the official currency of Spain and is used by 19 of the 27 European Union countries.
case 1 Retrieved evidence …The euro is the official currency of 20 of the 27 member states of the European Union.
Atomic Fact Chile is one of the countries with the highest forest cover in the world.
case 2 Retrieved evidence This is far more than the second greatest country in terms of forest area coverage, which is Russia . Russia has approximately ….
Long-form answer (Mintaka) The first book in the Hunger Games series, titled “The Hunger Games,” was published in 2008. (Total tokens: 16)
Atomic Fact The first book was published in 2008.
case 3 Retrieved evidence The Center for Fiction The Center for Fiction 2008 First Novel Prize …. This annual award was created in 2006 to honor the best first novel of the year.
Long-form answer (ExpertQA) A client with substance addiction, …. Secondly, Dialectical Behavior …. Additionally, Cognitive-Behavioral Therapy (CBT) …. Finally, a trauma-informed approach to addiction…. (Total tokens: 111)
Atomic Fact A client with substance addiction, … requires a comprehensive and integrative therapeutic approach.
case 4 Retrieved evidence … of childhood trauma in a range of settings that are particularly likely to serve this population like addictions, … in the context of the past trauma.
V-J Case Study
@@
When comparing the revised results of EFEC, RARR, and ARE in Table IX, it clearly highlights its advantages, as well as the shortcomings of RARR and EFEC. ARE performs minimal edits by focusing on atomic facts, successfully preserving the original intent while improving factuality and attribution. In contrast, RARR often makes substantial modifications, which can make it challenging to preserve the original intent of the content. EFEC performs even less satisfactorily. It makes significant changes to the original content, omitting entire sentences. Such drastic revisions can result in a final product that significantly deviates from the original content and intent.
As shown in Table X, cases 1 and 2 illustrate verification errors where fine-grained differences between evidence and facts are missed, leading to incorrect classifications (e.g., labeling “editing required” as “supportive” or vice versa). We randomly selected 50 samples for manual evaluation, finding a 23.1% error rate in misclassifying “editing required” as “irrelevant”. In case 1, the evidence cites 20 Eurozone countries while the fact mentions 19, but this discrepancy is not detected. Case 2 presents irrelevant evidence about Russia’s forests for a fact concerning Chile. Case 3 shows a decomposition failure where the LLM fails to extract the key subject “The Hunger Games” from a long-form answer, resulting in an invalid atomic fact. Case 4 highlights compounded failure in medical-domain content involving a long (111-token) multi-fact structure and inductive reasoning. The decomposition is insufficient, leading to ineffective retrieval and cascading errors across subsequent stages.
TABLE XI: The prompt templates used in ARE.
LLMs
Prompts
Long-form
answer
generation
You need to think step by step and answer my question.
Example 1:
[…]
Please answer the question and give the Explanation and
an answer.
Evidence
verifying
I will check some things you said. Here are some
examples for you to learn this process:
Example 1:
[…]
Now, please follow the above examples to inference.
Fact
expansion
Rewrite the fact into two short atomic phrases based on
Wikipedia. Ensure they contain more factual information
and can be easily supported by search engines.
Editing
You need use some evidences to check ’You said’ if
there are some difference and generate in ”My fix”.
Here are some examples to learn:
[…]
I will give you a new instance, please follow the above
example to fix a new one and start with ”My fix:”.
Backtracking
You are a linguist. Your task is to compose coherent
sentences using the provided atomic facts. While doing
so, refer to the structure of the given reference text —
including sentence count, subject order, and pronoun
usage — to maintain consistency and fluency.
Here are some examples:
{ “reference answer”: long-form answer
[{Molecular clause 1: [Atomic facts]},
{Molecular clause 2: [Atomic facts]},
{ Molecular clause 3: [Atomic facts] }]}
Now, think step by step to merge the given facts by
referring to the reference text. Generate the output in
coherent sentences starting with ”Sentences:”.
VI Conclusion